OpenStudy (lgbasallote):

evaluate \(\large \mathcal L \lbrace \sin kt \cos kt \rbrace\) with the aid of a trigonometric identity.

OpenStudy (turingtest):

\[\sin a\cos b=\frac12(\sin(a+b)+\sin(a-b))\]is probably the identity to use then it becomes like our earlier problem

OpenStudy (lgbasallote):

let me think this thorigh...

OpenStudy (lgbasallote):

through(

OpenStudy (lgbasallote):

through*

OpenStudy (turingtest):

thu*

OpenStudy (lgbasallote):

so im going to treat kt as a?

OpenStudy (turingtest):

k would be like a if you want to use de moivre theorem like @eliassaab did, but I would just split this sucker in two and integrate by parts like I did in the other

OpenStudy (lgbasallote):

so it's \[L \lbrace \frac{\sin (2kt)}{2} \rbrace\] ??

OpenStudy (lgbasallote):

well im not really familiar with that de moivre thingy

OpenStudy (turingtest):

me neither, that's why I'm trying to avoid it probably help here though I have not done a problem quite like this in deriving laplaces, so I will ponder it over my dinner :)

OpenStudy (turingtest):

While I'm eating I hope to watch it get solved, otherwise I'll try when I'm done :)

OpenStudy (lgbasallote):

hmm okay

OpenStudy (turingtest):

review example 4 for a basic idea of how to deal with the sine Laplace transform http://tutorial.math.lamar.edu/Classes/DE/LaplaceDefinition.aspx

OpenStudy (lgbasallote):

well i need to ask a few questions for checking so i guess i'll close this question for now :(

OpenStudy (anonymous):

Now as you did your other problem for Cos, you can do \[ \mathcal{L}_t[\sin (k t)](s)=\frac{k}{k^2+s^2}\\ Hence\\ \mathcal{L}_t[\sin (2 k t)](s)=\frac{2 k}{4 k^2+s^2}\\ \mathcal{L}_t\left[\frac{1}{2} \sin (2 k t)\right](s)=\frac{k}{4 k^2+s^2}\\ \] We are done since \[ \sin( k t )\cos(kt)=\frac{1}{2} \sin (2 k t) \]

OpenStudy (lgbasallote):

wow cool *_* so i was a bit on the track..

OpenStudy (lgbasallote):

so does this mean \[\mathcal L \lbrace \sin (nkt) \rbrace = \frac{nk}{s^2 + (nk)^2}\] @eliassaab ? because if yes i think i may be able to use it in the future

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

wow cool thanks